# Generalist Forecasting with Frozen Video Models via Latent Diffusion

### v1

@@ Abstract
Forecasting what will happen next is a critical skill for general-purpose systems that plan or act in the world at different levels of abstraction. In this paper, we identify a strong correlation between a vision model’s perceptual ability and its generalist forecasting performance over short time horizons. This trend holds across a diverse set of pretrained models—including those trained generatively—and across multiple levels of abstraction, from raw pixels to depth, point tracks, and object motion. The result is made possible by a novel generalist forecasting framework that operates on any frozen vision backbone: we train latent diffusion models to forecast future features in the frozen representation space, which are then decoded via lightweight, task-specific readouts. To enable consistent evaluation across tasks, we introduce distributional metrics that compare distributional properties directly in the space of downstream tasks and apply this framework to nine models and four tasks. Our results highlight the value of bridging representation learning and generative modeling for temporally grounded video understanding.
@@ 1 Introduction
The ability to see does not just reveal the present — it lets us anticipate the future and plan or act in the world accordingly. This capacity for visual forecasting is as critical for a gazelle dodging predators on the savanna as it is for a self-driving car navigating the urban jungle. At the same time, in most practical scenarios, the future is hard to predict. At any moment, countless possibilities lie ahead, and any model of the future must grapple with this inherent uncertainty.
@@
While modern computer vision models learn representations general enough to work across multiple levels of abstraction such as DINOv2 Oquab et al. (2023), most focus on generalist perception — tasks grounded in past and present frames with little to no stochasticity (Figure 1(a)). Evaluating vision models on these tasks tells us whether they understand what has already happened but may not reveal how well they can forecast what is to come.
@@
In this paper, we shift the focus from generalist video perception to generalist video forecasting—models that can predict future states of the world under uncertainty and across multiple levels of abstraction (Figure 1(b)), such that diverse perceptually relevant quantities can be decoded from a single predictive representation.
@@
We introduce a novel generalist forecasting framework built around a diffusion-based forecasting model, enabling forecasting across a range of frozen base video models and tasks. While recent work has explored using generative models for perception tasks (e.g.,Zhao et al. (2023); Luo et al. (2023); Li et al. (2023); Hedlin et al. (2023); Zhang et al. (2023); Bhattad et al. (2023); Xu et al. (2025)), the reverse—using perception models for forecasting—has been less explored. We fill this gap by showing that frozen video models trained for perception can be effectively repurposed for forecasting, and we establish a benchmark and strong baselines to support future explorations.
(a) Perception
(b) Forecasting
(c) Forecasting vs. perception performance
Figure 1: Forecasting performance strongly correlates with perceptual ability over short time horizons. While (a) perception is understanding current percepts, (b) forecasting predicts future states of the world. (c) We evaluate forecasting on pixels, point tracks, bounding box tracks, and depth using 10 samples per example and report the normalized max (min for lower-is-better metrics) performance per task. We compare with models’ performance on the perception-style variant of each task on observed frames. Data points correspond to model types. We find a linear correlation between forecasting and perception performance on these short-horizon context tasks.
Our diffusion-based forecasting framework enables direct, apples-to-apples comparison between perception- and synthesis-based models across all levels of visual abstraction. Our large-scale study reveals several key insights. First, forecasting ability is strongly correlated with perception performance—particularly for models pretrained with temporal video supervision as is clear from Figure 1(c). Second, video synthesis models like W.A.L.T. Gupta et al. (2024) meet or exceed the forecasting performance of models trained with mask-based objectives. Third, language supervision does not help forecasting performance and models trained solely on static images consistently underperform, highlighting the importance of temporal context in learning generalizable video representations.
We extend frozen, pretrained state-of-the-art video models to forecasting tasks in two stages. First, we fit a lightweight attention-based readout head to each model for each downstream task, following the perception-based paradigm of Carreira et al. (2024) (Figure 2(a)). This readout head maps from the space of frozen video representations to task outputs (e.g.  point tracks), trained using standard perception-based supervision. Then, we train a diffusion model to forecast future trajectories directly in the space of the frozen video representations (Figure 2(b)). During evaluation, we pass forecast trajectories through the readout head and assess their quality in the space of the downstream task (Figure 2(c)) using a suite of metrics that measure both realism and diversity—capturing the full dynamics and stochastic nature of the predicted futures.
Together, our findings and novel forecasting method outline a path toward video models that not only understand the present, but also reliably predict the future across the full spectrum of abstraction.
@@
Figure 2: Diffusion-based forecasting method from frozen vision model backbones. (a) Perception-style readouts: we train readout heads on frozen representations to perform downstream perception tasks like object detection on observed frames as in Carreira et al. (2024). We extend this setup to forecasting as follows. (b) Forecasting framework: We introduce a forecasting diffusion model that predicts future representations conditioned on frozen observed context representations. Pretrained readouts then decode these into future downstream abstractions, such as bounding boxes. (c) Forecasting across abstraction levels: We apply our approach to tasks spanning low to high-level structure—pixels, depth, point tracks, and object detections. Each example shows 4 observed frames and a sample from the 12 forecasted frames (frames 7, 10, 13, and 16). Our results show that frozen video representations can generalize to forecasting across a wide range of downstream tasks.
@@ 2 Related Work
Video (pixel) synthesis. Early video prediction models used recurrent architectures to directly model pixel intensities Ranzato et al. (2014); Oprea et al. (2022), but struggled with long-term dynamics. Probabilistic models like SVG-LP Denton and Fergus (2018) and GANs Clark et al. (2019); Tulyakov et al. (2018); Wang et al. (2020) improved visual quality by factoring content and motion. More recently, diffusion models Ho et al. (2020) have dominated video synthesis, with models like Sora OpenAI (2024), MovieGen Polyak et al. (2025), VideoPoet Kondratyuk et al. (2023) and W.A.L.T Gupta et al. (2024) leading the way. These models capture temporal dynamics through stochastic differential equations. While some diffusion-based works address video prediction Gu et al. (2023); Xing et al. (2024); Höppe et al. (2022); Ye and Bilodeau (2024); Yang et al. (2023), they are either language-guided or operate in implicit representation spaces. We leverage a diffusion model as a general forecasting engine across a range of visual abstractions.
@@
Task-specific video-based forecasting. Directly forecasting future pixels often fails to produce representations useful across abstraction levels. Luc et al. (2017) showed that forecasting semantic segmentation maps outperforms segmenting predicted RGB frames. They later proposed forecasting in the feature space of Mask R-CNN Luc et al. (2018), an approach similar to ours. Similarly,  Vondrick et al. (2016) forecast features of AlexNet to predict actions and objects in the future. However, we generalize this framework: rather than relying on task-specific networks, we forecast in the frozen representation space of large pretrained video models that support a broad range of downstream tasks.
@@
A separate line of work focuses on learning temporal dynamics from scratch using generative models or Neural Differential Equations (NDEs), such as Trajectory Flow Matching Zhang et al. (2024) and ImageFlowNet Liu et al. (2025). While these methods construct explicit, task-specific dynamical models, we ask how well general-purpose frozen video representations capture implicit dynamics that can be leveraged for forecasting via a separate diffusion-based module.
@@
Multi-task forecasting with frozen video representations. Instead of designing task-specific forecasting models, several works have explored forecasting directly in the frozen representation space of large pretrained video models, leveraging their generality across downstream tasks. In this setup, future representations are predicted by learning lightweight forecasting heads on top of frozen features. Variants of this approach appear in recent work Rajasegaran et al. (2025); Karypidis et al. (2024). Rajasegaran et al. (2025) pretrain autoregressive next-token prediction models on large-scale video data and evaluate the resulting frozen features using probing tasks such as short-term interaction anticipation.
@@
Most related to our work, DINO-Foresight Karypidis et al. (2024) forecasts frozen DINOv2 Oquab et al. (2023) features using a masked transformer, evaluating downstream tasks such as segmentation, depth, and surface normals at a single future time point. Unlike our approach, DINO-Foresight assumes deterministic futures and performs single-step prediction, while we model uncertainty and full temporal trajectories.
@@ 3 Method
While prior works have explored forecasting directly in pixel space, we hypothesize that latent spaces should be better because they make the scene structure more clear and remove non-semantic, hard-to-predict details, which should make prediction easier. Therefore, we start with frozen pretrained models, and use them to both represent the conditioning (past) video and the future video that we wish to predict. We develop a two-stage forecasting framework built around a diffusion-based forecasting module that operates directly in the space of frozen video representations. This setup allows us to extend representations—trained for perception or pixel synthesis—to forecasting tasks without fine-tuning. We first train lightweight readout heads to decode task-specific outputs from frozen representations. Then, we train a diffusion model to forecast future latent trajectories in the space of the frozen video representations. These forecast representations are passed through the same readouts, enabling the evaluation of nondeterministic futures across multiple semantic levels. The full pipeline is illustrated in Figure 2.
@@ 3.1 Latent Forecasting via Diffusion
We forecast future representations using a conditional denoising diffusion model Ho et al. (2020). Given a sequence of frozen representations up to time $t$, the diffusion model generates future latent trajectories for times $t+1...T$ conditioned on the past (Figure 2(b)). This design allows us to model the multi-modal nature of the future and capture stochastic temporal dynamics. Unlike pixel-space synthesis, our model operates in the latent space of each frozen backbone, making it architecture-agnostic and capable of comparing perception and synthesis models under the same framework. We train one diffusion forecasting module per frozen video model under consideration. In our experiments, we condition the diffusion model on the encodings of 4 past frames and model the latents of the 12 next frames. The diffusion models the latent encodings of these future frames jointly in time.
@@ 3.2 Task Readout Heads
We decode the sampled latent trajectories using the previously trained readout heads to evaluate forecast futures. This allows us to assess prediction quality across different abstraction levels using task-appropriate metrics (Figure 2(c)). We train a lightweight attention-based readout head to decode the output of each frozen model into a task-specific prediction space. We focus on four tasks that vary in their level of abstraction. These are pixels, depth, point tracks, and bounding boxes. The readout heads for the four tasks follow those of Carreira et al. (2024) with the architecture of the depth readout head also used (but trained separately) for the pixel readout task (Figure 2(a)). These readouts are trained using standard supervised losses and provide a shared interface for comparing models across different architectures and pretraining paradigms. Importantly, readout heads are trained only on observed (past and present) frames and remain fixed during forecasting. We train a readout head for each frozen video model and downstream task pair.
@@ 3.3 Evaluation Metrics
Per Example Metrics. For each example, we take 10 samples from our diffusion model and report the statistics of task-specific metrics from the ground truth over each sample. For pixel prediction, we use PSNR. For depth prediction, mean absolute relative error. For point tracks, we use Jaccard Distance. For box tracking, we use intersection over union. In order to account for the stochastic nature of forecasting, we report mean, minimum, and maximum of per-example samples for the given metric.
@@
Fréchet Distance. In order to capture the inherent stochasticity of forecasting the future, we must allow for a distribution over possible future trajectories and ensure that this forecasted distribution is similar to that of the target ground truth data. We therefore compute the Fréchet Distance (FD) Fréchet (1957), a distribution distance metric comparing the forecasted versus the ground truth set distribution over trajectories, in the output representation of each task. While Ng et al. (2022) employed FD for motion forecasting evaluation and Thakkar et al. (2025) for self-driving cars, action, and object interactions, we extend the use of FD for evaluating forecasts of pixels, depth, point tracks, and object bounding box tracks. Specifically, we first represent forecasts as points in some fixed dimensional space. For point tracks and box tracks, we represent each trajectory as, respectively, a vector in a 24-dimensional (2D coordinates over 12 future frames) and 48-dimensional (4 coordinates over 12 frames) spaces. For depth and pixels, we downsample each of the 12 output frames to $14\times 14$ patches, leading to a 2352-dimensional representation space. We then fit multivariate Gaussians to the predicted and ground truth distributions in this space, and compute the Fréchet distance Dowson and Landau (1982) between them. To ensure our output is always of a fixed size, we filter out trajectories that do not contain all the available data points (e.g. point tracks that are not visible across all target frames due to occlusion). Note that while Heusel et al. (2017); Unterthiner et al. (2019) compute FD in the Inception embedding space, there is no Inception here. We compute FD directly in the output representation of each downstream task.
@@
Variance. While FD considers the realism and stochasticity of the forecasted futures, it is prudent to pair it with a measure of the variance over these futures to ensure that the forecasted futures are as diverse as the ground truth ones. We report the variance of the trajectories over the temporal axis, averaged over all other dimensions. This specifically assesses whether methods always forecast static future trajectories, which may be realistic but certainly not diverse.
Together, our modular framework decouples representation learning, task decoding, and forecasting, enabling fair and extensible evaluation across a wide range of pretrained video models. By using a diffusion model in representation space, we unify evaluation across perception and synthesis-based video models at multiple abstraction levels.

### v2

# Frozen Forecasting: A Unified Evaluation

@@ Abstract
Forecasting future events is a fundamental capability for general-purpose systems that plan or act across different levels of abstraction. Yet, evaluating whether a forecast is “correct” remains challenging due to the inherent uncertainty of the future. We propose a unified evaluation framework for assessing the forecasting capabilities of frozen vision backbones across diverse tasks and abstraction levels. Rather than focusing on single time steps, our framework evaluates entire trajectories and incorporates distributional metrics that better capture the multimodal nature of future outcomes. Given a frozen vision model, we train latent diffusion models to forecast future features directly in its representation space, which are then decoded via lightweight, task-specific readouts. This enables consistent evaluation across a suite of diverse tasks while isolating the forecasting capacity of the backbone itself. We apply our framework to nine diverse vision models, spanning image and video pretraining, contrastive and generative objectives, and with or without language supervision, and evaluate them on four forecasting tasks, from low-level pixel predictions to high-level object motion. We find that forecasting performance strongly correlates with perceptual quality and that the forecasting abilities of video synthesis models are comparable or exceed those pretrained in masking regimes across all levels of abstraction. However, language supervision does not consistently improve forecasting. Notably, video-pretrained models consistently outperform image-based ones.
@@ 1 Introduction
The ability to see does not just reveal the present. It lets us anticipate the future and plan or act in the world accordingly. This capacity for visual forecasting is as critical for a gazelle dodging predators on the savanna as it is for a self-driving car navigating the urban jungle. At the same time, in most practical scenarios, the future is hard to predict. At any moment, countless possibilities lie ahead, and any model of the future must grapple with this inherent uncertainty.
@@
While modern computer vision models learn representations general enough to work across multiple levels of abstraction such as DINOv2 [oquab2023dinov2] and 4DS [carreira2024scaling], most focus on perception — tasks grounded in past and present frames with little to no stochasticity. Several methods of evaluations have been developed for self-supervised perception tasks, such as linear readouts, nearest-neighbors, cross-attention based, etc. However, evaluating vision models on these tasks tells us whether they understand what has already happened but may not reveal how well they can forecast what is to come.
@@
In this paper, we shift the focus from evaluating the video perception capabilities of vision models to evaluating their video forecasting capabilities—assessing learned visual representations that can be used to predict future states of the world under uncertainty and across multiple levels of abstraction, such that diverse perceptually relevant quantities can be decoded from a single predictive representation.
@@
We propose a unified forecasting evaluation framework built around a diffusion-based forecasting model, enabling forecasting across a range of frozen base video models and tasks: pixels, depth, point tracks, and bounding boxes. While recent work has explored using generative models for perception tasks (e.g.,[zhao2023unleashing, luo2023diffusion, li2023your, hedlin2023unsupervised, zhang2023tale, bhattad2023stylegan, xu2024matters]), the reverse—using perception models for forecasting—has been less explored. We fill this gap by showing that frozen video models trained for perception can be effectively repurposed for forecasting, and we establish a benchmark and strong baselines to support future explorations.
@@
Figure 1: Diffusion-based forecasting evaluation framework of frozen vision model backbones. (a) Perception-style readouts: we train readout heads on frozen representations to perform downstream perception tasks like object detection on observed frames as in [carreira2024scaling]. We extend this setup to forecasting as follows. (b) Forecasting framework: We introduce a forecasting diffusion model that predicts future representations conditioned on frozen observed context representations. Pretrained readouts then decode these into future downstream abstractions, such as bounding boxes. (c) Forecasting across abstraction levels: We apply our approach to evaluate forecasting on tasks spanning low to high-level structure—pixels, depth, point tracks, and object detections. Each example shows 4 observed frames and a sample from the 12 forecast frames (frames 7, 10, 13, and 16). Our results show that frozen video representations can generalize to forecasting across a wide range of downstream tasks.
We extend frozen, pretrained state-of-the-art video models to forecasting tasks in two stages. First, we fit a lightweight attention-based readout head to each model for each downstream task, following the perception-based paradigm of [carreira2024scaling] (Figure 1(a)). This readout head maps from the space of frozen video representations to task outputs (e.g.  point tracks), trained using standard perception-based supervision. Then, we train a diffusion model to forecast future trajectories directly in the space of the frozen video representations (Figure 1(b)). During evaluation, we pass forecast trajectories through the readout head and assess their quality in the space of the downstream task (Figure 1(c)) using a suite of metrics that measure both realism and diversity—capturing the full dynamics and stochastic nature of the predicted futures.
Our diffusion-based forecasting framework enables direct, apples-to-apples comparison between perception- and synthesis-based models across all levels of visual abstraction. Our large-scale study reveals several key insights. First, forecasting ability is generally correlated with perception performance, but this association starts to break down with the strongest models. Second, video synthesis models like WALT [gupta2024walt] meet or exceed the forecasting performance of similarly-sized models trained with mask-based objectives when evaluated with a distribution-sensitive metric. Third, language supervision does not help forecasting performance. Fourth, models trained solely on static images consistently underperform, highlighting the importance of temporal context in learning generalizable video representations.
@@ 2 Related Work
Video (pixel) synthesis. Early video prediction models used recurrent architectures to directly model pixel intensities [ranzato2014video, 9294028], but struggled with long-term dynamics. Probabilistic models like SVG-LP [denton2018stochastic] and GANs [clark2019adversarial, tulyakov2018mocogan, wang2020imaginator] improved visual quality by factoring content and motion. More recently, diffusion models [ho2020denoising] have dominated video synthesis, with models like Sora [openai-2024], MovieGen [polyak2025moviegencastmedia], VideoPoet [kondratyuk2023videopoet] and WALT [gupta2024walt] leading the way. These models capture temporal dynamics through stochastic differential equations. While some diffusion-based works address video prediction [gu2023seer, xing2024aid, hoppe2022diffusion, ye2024stdiff, yang2023video], they are either language-guided or operate in implicit representation spaces. We leverage a diffusion model as a general forecasting engine across a range of visual abstractions. In addition, there has been work on diffusion video models which are autoregressive in time.  [hu2024acdit] is autoregressive on spatial-temporal blocks and leverages diffusion to generate unseen blocks.  [zhang2025generative] and  [huang2025self] are autoregressive on past video frames and use diffusion to forecast the latents of future frames, similar to our proposed architecture. These approaches to autoregressive diffusion are potential design alternatives to the forecasting module in our evaluation framework that would be interesting to explore in future work.
@@
Task-specific video-based forecasting. Directly forecasting future pixels often fails to produce representations useful across abstraction levels. [luc2017predicting] showed that forecasting semantic segmentation maps outperforms segmenting predicted RGB frames. They later proposed forecasting in the feature space of Mask R-CNN [luc2018predicting], an approach similar to ours. Similarly,  [vondrick2016representation] forecast features of AlexNet to predict actions and objects in the future. Similarly, other previous works [saric2020warp, lin2021predictive] also focus specifically on the task of forecasting segments or pixel interpolation [argaw2022long]. However, we generalize this framework: rather than relying on task-specific networks, we forecast in the frozen representation space of large pretrained video models that support a broad range of downstream tasks.
@@
A separate line of work focuses on learning temporal dynamics from scratch using generative models or Neural Differential Equations (NDEs), such as Trajectory Flow Matching [zhang2024trajectory] and ImageFlowNet [liu2025imageflownet]. While these methods construct explicit, task-specific dynamical models, we ask how well general-purpose frozen video representations capture implicit dynamics that can be leveraged for forecasting via a separate diffusion-based module.
@@
Multi-task forecasting with frozen video representations. Instead of designing task-specific forecasting models, several works have explored forecasting directly in the frozen representation space of large pretrained video models, leveraging their generality across downstream tasks. In this setup, future representations are predicted by learning lightweight forecasting heads on top of frozen features. Variants of this approach appear in recent work [rajasegaran2025empirical, karypidis2024dinoforesightlookingfuturedino]. [rajasegaran2025empirical] pretrain autoregressive models on large-scale video data and evaluate the resulting frozen features using probing tasks such as short-term interaction anticipation.
@@
Most related to our work, DINO-Foresight [karypidis2024dinoforesightlookingfuturedino] and Back to the Features [baldassarre2025back] forecast frozen DINOv2 [oquab2023dinov2] features using a masked transformer and autoregressive model, respectively, evaluating downstream tasks such as segmentation, depth, and surface normals at a single future time point. Unlike our approach, these works assume deterministic futures and perform single-step prediction, while we model uncertainty and evaluate the entire distribution of future trajectories.
Stochastic approaches to forecasting. In many contexts, visual forecasting is an inherently stochastic problem. A simple deterministic regressor may not necessary capture the full diversity of possible future outcomes. Some previous approaches have attempted to address this stochasticity. For example, [bhattacharyya2019bayesian] proposes a Bayesian model that jointly captures epistemic and observation aleatoric of future states.  [makansi2020multimodal] uses mixture density networks to estimate the location of objects like pedestrians and vehicles from an egocentric view. In this paper, we use a diffusion model to directly learn the continuous distribution of future features.
@@ 3 Method
While prior works have explored forecasting directly in pixel space, we hypothesize that latent spaces should be better because they make the scene structure more clear and remove non-semantic, hard-to-predict details, which should make prediction easier. Therefore, we start with frozen pretrained models, and use them to both represent the conditioning (past) video and the future video that we wish to predict. We develop a two-stage forecasting evaluation framework built around a diffusion-based forecasting module that operates directly in the space of frozen video representations. This setup allows us to extend representations trained for perception or pixel synthesis to forecasting tasks without fine-tuning. We first train lightweight readout heads to decode task-specific outputs from frozen representations. Then, we train a diffusion model to forecast future latent trajectories in the space of the frozen video representations. These forecast representations are passed through the same readouts, enabling the evaluation of nondeterministic futures across multiple semantic levels. The full pipeline is illustrated in Figure 1.
@@ 3.1 Latent Forecasting via Diffusion
We forecast future representations using a conditional denoising diffusion model [ho2020denoising]. Given a sequence of frozen representations up to time $t$, the diffusion model generates future latent trajectories for times $t+1\ldots T$ conditioned on the past (Figure 1(b)). Unlike pixel-space synthesis, our model operates in the latent space of each frozen backbone, making it architecture-agnostic and capable of comparing perception and synthesis models under the same framework.
We train one diffusion forecasting module per frozen video model under consideration. In our experiments, we condition the diffusion model on the latent encodings of $t=4$ past frames after applying layer normalization, as it was found that the diffusion model would occasionally struggle to forecast in unnormalized latent space. We model latent encodings of a $T=16$ frame clip, which includes the 4 past frames and 12 future frames ($16\times 224\times 224\times 3$ clip). The diffusion models the latent encodings of all these frames jointly in time. Even though the diffusion model may already have information on the first $t$ frames for conditioning, the entire clip is modeled jointly to account for temporally-entangled features.
The diffusion forecasting model never takes as conditioning any direct pixel information, only latent encodings from a given video model. This pipeline is the same whether the encoder is a video or image model. In the image encoder case, the latents are the stacked result of the image encoder on all frames.
@@ 3.2 Task Readout Heads
We decode the sampled latent trajectories using the previously trained readout heads to evaluate forecast futures. This allows us to assess prediction quality across different abstraction levels using task-appropriate metrics (Figure 1(c)). We train a lightweight attention-based readout head to decode the output of each frozen model into a task-specific prediction space. We focus on four tasks that vary in their level of abstraction. These are pixels, depth, point tracks, and bounding boxes. The readout heads for the four tasks follow those of [carreira2024scaling] with the architecture of the depth readout head also used (but trained separately) for the pixel readout task (Figure 1(a)). These readouts are trained using standard supervised losses and provide a shared interface for comparing models across different architectures and pretraining paradigms. Importantly, readout heads are trained only on observed (past and present) frames and remain fixed during forecasting. We train a readout head for each frozen video model and downstream task pair. During inference, the transformer-based readout head processes features from all frames, conditional and forecast, simultaneously via full attention. The loss from the readout heads is not backpropagated to the diffusion model.
@@ 3.3 Evaluation Metrics
Per Example Metrics. For each example, we take 10 samples from our diffusion model and report the statistics of task-specific metrics from the ground truth over each sample. For pixel prediction, we use PSNR. For depth prediction, mean absolute relative error. For point tracks, we use Jaccard Distance. For box tracking, we use intersection over union. In order to account for the stochastic nature of forecasting, we report mean, minimum, and maximum of per-example samples for the given metric. For relative comparison, we also report the perception, or standalone performance on the ground truth latents on all frames, of the readout heads alongside the per example metrics.
@@
Fréchet Distance. In order to capture the inherent stochasticity of forecasting the future, we must allow for a distribution over possible future trajectories and ensure that this forecast distribution is similar to that of the target ground truth data. We therefore compute the Fréchet Distance (FD) [frechet1957distance], a distribution distance metric comparing the forecast versus the ground truth set distribution over trajectories, in the output representation of each task. While [ng2022learning] employed FD for motion forecasting evaluation and [thakkar2025poly] for self-driving cars, action, and object interactions, we extend the use of FD for evaluating forecasts of pixels, depth, point tracks, and object bounding box tracks. Specifically, we first represent forecasts as points in some fixed dimensional space. For point tracks and box tracks, we represent each trajectory as, respectively, a vector in a 24-dimensional (2D coordinates over 12 future frames) and 48-dimensional (4 coordinates over 12 frames) spaces. For depth and pixels, we downsample each of the 12 output frames to $14\times 14$ patches, leading to a 2352-dimensional representation space. We then fit multivariate Gaussians to the predicted and ground truth distributions in this space, and compute the Fréchet distance [dowson1982frechet] between them. To ensure our output is always of a fixed size, we filter out trajectories that do not contain all the available data points (e.g. point tracks that are not visible across all target frames due to occlusion). Note that while [heusel2017gans, unterthiner2019fvd] compute FD in the Inception embedding space, there is no Inception here. We compute FD directly in the output representation of each downstream task. Explicit details are provided in Appendix 0.A.1.
@@
Variance. While FD considers the realism and stochasticity of the forecast futures, it is prudent to pair it with a measure of the variance over these futures to ensure that the forecast futures are as diverse as the ground truth ones. We report the variance of the trajectories over the temporal axis, averaged over all other dimensions. This specifically assesses whether methods always forecast static future trajectories, which may be realistic but certainly not diverse.